Nitrogen tetroxide dissociates as follows: N2O4 --> 2NO2
If a 100ml flask contains 1.00gm of N2O4 that is 30% dissociated at 313K, what is the pressure inside the flask in atmospheres?
1
Expert's answer
2016-12-09T03:12:11-0500
First we have to calculate an average molar mass of the equilibrium mixture M(N2O4/NO2) = 92 g/mol*0.7 + 46 g/mol*0.3 = 78 g/mol.
Now, according to the gas state equation P = m*R*T/M/V = 10-6 kg * 8.31 J/mol K * 313 K / 78*10-3 kg/mol * 0.1*10-3 m3 = 333 Pa.